The molar mass of carbon dioxide gas is 44.0 g/mol. How many molecules are in 52.0 g of carbon dioxide gas?
1
Expert's answer
2020-11-11T10:35:53-0500

We use Avogadro's number of "6.022\times 10^{23}" molecules per mole to calculate the number of molecules from the number of moles represented by "52.0g" of carbon dioxide.

The molecular weight of "CO_2" "=44.0g\/mol"

"=""52.0g\over 44.0g/mol" "\times" "6.022\times 10^{23} molecules\over1mol"

"=7.117\times 10^{23} molecules of CO_2"